# Fostering Human Learning in Sequential Decision-Making: Understanding the Role of Evaluative Feedback

### v3

@@ 1 Introduction
The rest of the manuscript is structured as follows. Sec. 2 presents background and problem formulation, and includes a discussion of the ToH structure, the application of maximum entropy IRL for learning human rewards, and the development of computational models aimed at integrating evaluative feedback into human decision-making processes. In Sec. 3, we provide details of the ToH experiments, conducted through the Amazon Mechanical Turk (AMT) platform, alongside the discussion of the various evaluative feedback strategies employed during these experiments. We discuss and analyze the experiment results in Section 4 and finally conclude in Sec 5.
@@ 2 Background and Problem Formulation
Fig 1: State space of a $4$-disk ToH with $81$ states. Each state corresponds to a unique configuration of the disks on three pegs and edges encode allowed transitions between states. The task is to reach the configuration associated with a randomly selected target state (for example $2201$ in this figure). Warmer colors are associated with the higher value function (see Sec. 2.1 for discussion).
@@ 2.3 Modeling human sequential decision-making under feedback
  • Model $1$ Ignore feedback: This baseline model operates under the assumption that evaluative feedback isn’t directly integrated into human decision-making processes. Instead, individuals are postulated to focus on maximizing the long-term value derived from their personal reward functions. In this framework, evaluative feedback plays an indirect role by shaping and refining these reward functions. This is the default model studied in Sec. 2.2. The model encompasses $|f|$ learned parameters.
@@
We investigate these models in Sec 4.3 to understand how humans incorporate evaluative feedback in their decision-making.
@@ 4.3 Modeling Human Decision-Making under evaluative feedback
In Sec. 4.1 and 4.2, we have demonstrated the pivotal role of evaluative feedback in enhancing learning and performance within the context of the ToH puzzle. In this section, we delve into exploring models that aim to elucidate the mechanisms through which humans integrate evaluative feedback into their decision-making processes.
In our analysis, we explore four distinct models for incorporating feedback into human decision-making, as detailed in Sec. 2.3. For each of these models, we calculate both the Akaike information criterion (AIC) [40] and Bayesian information criterion (BIC) [41] to identify the most suitable model. For a given model, the AIC and BIC are defined as:
@@
Upon completing this partitioning process for the $200$ trajectories obtained from the training tasks, we arrived at six groups, each containing a respective number of trajectories: $41,34,27,31,32,35$. Within each group, we subject all four models to testing, making appropriate modifications to either the Q-function or the reward function as discussed in Sec. 2.3. To estimate the unknown parameters for each model, we employ maximum entropy IRL, optimizing the $\log$-likelihood as defined in (6) while applying an $\mathcal{L}_{1}$ penalty.
@@
Similar to Sec. 4.2, we investigate two settings for the predefined features: one that allows non-zero rewards in all states and another that restricts rewards to a subset of $8$ states. Table 2 presents the AIC and BIC values (normalized by the number of observations) for different models within each group when non-zero rewards are permitted for all $81$ states. It’s worth noting that, while Model $2$, where $Q^{\prime}(s,a)=Q(s,a)+k\hat{H}(s,a)$, emerges as the best fit according to AIC for the majority of the groups, Model $4$, where $Q^{\prime}(s,a)=k\hat{H}(s,a)$, is selected as the best fit under the BIC criterion. This preference for Model $4$ under BIC is attributed to the significant difference in the number of learned parameters between the two models, with BIC favoring the model with fewer learned parameters. Indeed, when considering non-zero rewards for all $81$ states, it leads to a preference for the model with just a single learned parameter.

### v4

@@ 1 Introduction
The rest of the manuscript is structured as follows. Section 2 presents background and problem formulation, and includes a discussion of the ToH structure, the application of maximum entropy IRL for learning human rewards, and the development of computational models aimed at integrating evaluative feedback into human decision-making processes. In Section 3, we provide details of the ToH experiments, conducted through the Amazon Mechanical Turk (AMT) platform, alongside the discussion of the various evaluative feedback strategies employed during these experiments. We discuss and analyze the experiment results in Section 4 and finally conclude in Section 5.
@@ 2 Background and Problem Formulation
Fig 1: State space of a $4$-disk ToH with $81$ states. Each state corresponds to a unique configuration of the disks on three pegs and edges encode allowed transitions between states. The task is to reach the configuration associated with a randomly selected target state (for example $2201$ in this figure). Warmer colors are associated with the higher value function (see Section 2.1 for discussion).
@@ 2.3 Modeling human sequential decision-making under feedback
  • Model $1$ Ignore feedback: This baseline model operates under the assumption that evaluative feedback isn’t directly integrated into human decision-making processes. Instead, individuals are postulated to focus on maximizing the long-term value derived from their personal reward functions. In this framework, evaluative feedback plays an indirect role by shaping and refining these reward functions. This is the default model studied in Section 2.2. The model encompasses $|f|$ learned parameters.
@@
We investigate these models in Section 4.3 to understand how humans incorporate evaluative feedback in their decision-making.
@@ 4.3 Modeling Human Decision-Making under evaluative feedback
In Section 4.1 and 4.2, we have demonstrated the pivotal role of evaluative feedback in enhancing learning and performance within the context of the ToH puzzle. In this section, we delve into exploring models that aim to elucidate the mechanisms through which humans integrate evaluative feedback into their decision-making processes.
In our analysis, we explore four distinct models for incorporating feedback into human decision-making, as detailed in Section 2.3. For each of these models, we calculate both the Akaike information criterion (AIC) [40] and Bayesian information criterion (BIC) [41] to identify the most suitable model. For a given model, the AIC and BIC are defined as:
@@
Upon completing this partitioning process for the $200$ trajectories obtained from the training tasks, we arrived at six groups, each containing a respective number of trajectories: $41,34,27,31,32,35$. Within each group, we subject all four models to testing, making appropriate modifications to either the Q-function or the reward function as discussed in Section 2.3. To estimate the unknown parameters for each model, we employ maximum entropy IRL, optimizing the $\log$-likelihood as defined in (6) while applying an $\mathcal{L}_{1}$ penalty.
@@
Similar to Section 4.2, we investigate two settings for the predefined features: one that allows non-zero rewards in all states and another that restricts rewards to a subset of $8$ states. Table 2 presents the AIC and BIC values (normalized by the number of observations) for different models within each group when non-zero rewards are permitted for all $81$ states. It’s worth noting that, while Model $2$, where $Q^{\prime}(s,a)=Q(s,a)+k\hat{H}(s,a)$, emerges as the best fit according to AIC for the majority of the groups, Model $4$, where $Q^{\prime}(s,a)=k\hat{H}(s,a)$, is selected as the best fit under the BIC criterion. This preference for Model $4$ under BIC is attributed to the significant difference in the number of learned parameters between the two models, with BIC favoring the model with fewer learned parameters. Indeed, when considering non-zero rewards for all $81$ states, it leads to a preference for the model with just a single learned parameter.